# Wildfire Risk Prediction: A Review

### v3

@@ 4.1.3 Fuel types
Another method to describe fuel properties, aside from fuel moisture code and fuel load, is by categorizing fuels into homogeneous groups that exhibit similar characteristics under specified burning conditions. These groups are defined as "fuel types" or "fuel models" (Arroyo et al., 2008). The description of fuel types is generally complex, taking into account various fuel characteristics such as vegetation type, canopy height, canopy base height, canopy bulk density, vegetation cover ratio, horizontal and vertical continuity, fuel load, fuel moisture, and biomass (Chuvieco et al., 2003; Pettinari and Chuvieco, 2020; Aragoneses and Chuvieco, 2021). In recent years, various standardized fuel classification systems have been proposed in different regions (Fig.1), including the NFDRS (Cohen, 1985), Northern Forest Fire Laboratory (NFFL) (Albini, 1976; Burgan, 1984), Fuel Characteristics Classification System (FCCS) (Sandberg et al., 2001; Ottmar et al., 2007), McArthur fuel types (McArthur, 1966, 1967), Canadian Fire Behaviour Prediction (FBP) (Group et al., 1992), and Prometheus (Riaño et al., 2002). These systems are either fire rating systems or fuel classification systems defined by wildfire simulation methods, or simplified versions of such systems (Arroyo et al., 2008).
@@
It is noteworthy that various vegetation state description parameters are generally used in studies evaluating their effectiveness in wildfire risk modeling but are rarely directly utilized in wildfire risk assessment models. Commonly used wildfire risk assessment models often directly employ raw multispectral or microwave data or indices derived from these data. This could be due to several reasons:
@@
Lastly, data-driven algorithms are becoming increasingly popular, and these algorithms can either rely on or bypass manual feature engineering, allowing them to directly learn the necessary abstractions from features or raw data.
4.1.4 Fuel continuity
CBH is defined as the height above which there is sufficient canopy fuel per unit volume to support vertical fire spread into the canopy (Scott, 2001). It has widespread application in various wildfire behavior models (Hall and Burke, 2006). CBH can be estimated using allometric equations based on species, diameter at breast height (DBH), and tree height, or it can be retrieved from remote sensing data. However, due to the limitations of passive optical remote sensing signals and synthetic aperture radar (SAR) in penetrating the canopy, studies using multi-angle observations or bidirectional reflectance distribution function (BRDF) retrievals to obtain three-dimensional vegetation structure information are rare in wildfire prediction (Pisek et al., 2015; Canisius and Chen, 2007). More commonly, LiDAR is used to estimate CBH at the individual tree or local scale (Andersen et al., 2005; Luo et al., 2018; Wang et al., 2022). Some studies suggest that while the definition of CBH is clear, its practical application is complicated by the lack of a clear fuel threshold per unit volume and the potential neglect of fine fuels, which can lead to errors in modeling canopy fire risk (Cruz et al., 2004; Kramer et al., 2014). As a result, a quick and simple method to estimate ladder fuels is by estimating the coverage area or coverage rate of low-lying fuels as a representation of ladder fuel (Clark et al., 2009; Wing et al., 2012).
@@ 4.2 Weather and climate conditions
Weather and climate factors are major drivers of wildfire activity, influencing the occurrence, burned area, and fire behavior (Swetnam and Betancourt, 1990; Bessie and Johnson, 1995; Abatzoglou and Kolden, 2013). Firstly, weather conditions are central to almost all global fire danger risk systems. High temperatures, low relative humidity, insufficient precipitation, and strong winds have been proven to be key factors in the ignition of fires in the short term (Keane et al., 2001; Zacharakis and Tsihrintzis, 2023b), and they also affect fire size and spread (Abatzoglou and Kolden, 2011). On the other hand, climate conditions generally refer to the long-term average weather conditions of a region and their variations. Some studies specifically define these as the collective atmospheric conditions months or even years before the start of the fire season, emphasizing the use of prior climate conditions to convert fuel conditions into wildfire potential (Abatzoglou and Kolden, 2013). This approach enables predictions of large-scale fire risks over seasonal, annual, or even multi-decadal time scales.
@@ 4.2.1 Weather conditions
Short-term weather conditions focus primarily on predicting the impact of weather conditions over quarters, months, weeks, or days on fire risk. These include short-term temperature, wind, relative humidity, soil moisture, precipitation, and lightning, as well as corresponding indices like the Palmer Drought Severity Index (PDSI) and the Fire Weather Index (FWI). Temperature and various humidity conditions affect fuel conditions, such as evapotranspiration and moisture content, which in turn influence ambient humidity conditions and wildfire risk. Wind direction, wind speed, and other related metrics like Mean Wind Power Density (MWPD) and Mean Wind Speed (MWS) affect fire spread and intensity, making them commonly used parameters in wildfire prediction. Unlike other parameters, lightning can act as an ignition factor for wildfires.
Short-term weather parameters are widely used in wildfire risk prediction. For example, Tavakkoli Piralilou et al. (2022) used temperature, precipitation, MWPD, and MWS as parameters in their wildfire prediction model for the forested areas of Gilan Province, Iran. In another study conducted in Pakistan, Kanwal et al. (2023) utilized 16 meteorological parameters, including monthly average precipitation, evapotranspiration, wind speed, soil temperature, humidity, heat flux, albedo, average land surface temperature, soil moisture, actual evapotranspiration, moisture deficit, downward surface shortwave radiation, precipitation accumulation, minimum/maximum temperature, and vapor pressure. An assessment of their importance indicated that heat flux, evapotranspiration, and vapor pressure were relatively more important compared to other meteorological parameters, although the differences in importance were not significant. Bergado et al. (2021) used seven meteorological factors in their wildfire prediction study for Victoria, Australia: daily maximum temperature, daily minimum relative humidity, daily total solar radiation, daily total precipitation, daily average wind speed, daily average wind direction, and lightning frequency. The lightning data was derived from an annual lightning climatology dataset that provided estimates of lightning density for each day of the year. Their study found that total precipitation, lightning density, and surface temperature consistently held high weight in all models. In contrast, Quan et al. (2023) used various meteorological variables from the ERA5 dataset, including precipitation, relative humidity, temperature, wind speed, and the Keetch-Byram Drought Index, to predict wildfire risk in parts of the western Tibetan Plateau in China. The study revealed that relative humidity was the most critical factor in the occurrence of forest and grassland fires in the region. Higher relative humidity leads to higher Dead Fuel Moisture Content (DFMC), thereby reducing the likelihood of ignition.
@@ 4.2.2 Climate conditions
Climate conditions mainly describe annual or multi-year meteorological parameters such as temperature, humidity, precipitation, and evaporation. Unlike meteorological conditions that influence short-term fuel status, climate conditions can control fuel types, distribution, life cycles, and more, thereby influencing fuel load and fuel continuity (Aldersley et al., 2011; Jaafari et al., 2017; Tavakkoli Piralilou et al., 2022). However, the distinction between climate and meteorological conditions lies in that climate variables are often the primary drivers of large-scale wildfire activity (Pollina et al., 2013; Jain et al., 2020). This is because fuel characteristics respond to climatic accumulation before the start of the fire season (Chen et al., 2016). Therefore, climate conditions are more suitable for predicting large-scale and long-term wildfires. Climate conditions can also explain strong seasonal variations in wildfire risk in certain regions (Martell et al., 1989), especially considering the intensification of climate change in recent years, which has raised concerns about its impact on the occurrence of large-scale wildfires over the long term (Prior and Eriksen, 2013; Jolly et al., 2015; Jain et al., 2024).
@@
Specifically, Chen et al. (2016) used large-scale sea surface temperature anomalies to predict the severity of the fire season in South America on a seasonal time scale. Their study found that using a single ocean climate index could predict about 48% of the burned area globally for up to three months or longer before the peak burning month. In a study by Jaafari et al. (2019) in the Zagros ecoregion of western Iran, they used three climate conditions: annual temperature, rainfall, and wind effects. However, the results showed that only the annual temperature parameter was relatively important. Similarly, Naderpour et al. (2020) studied the northern beaches of New South Wales, Australia, using three climate factors: annual rainfall, annual humidity, and annual wind speed. Holdrege et al. (2024) studied the Great Basin sagebrush regions across 13 western U.S. states and used three climate factors: annual mean temperature, annual precipitation, and the proportion of summer (June to August) precipitation to total annual precipitation. The study found that areas with low summer precipitation proportion, medium to high annual precipitation, and high temperatures had the highest wildfire probability. Similarly, Aldersley et al. (2011) used annual cumulative precipitation, annual lightning frequency, monthly precipitation frequency, monthly average temperature, and climate data from the British Atmospheric Data Centre (BADC) to study the drivers of monthly fire area on a global and regional scale. They found that climate factors such as high temperatures, moderate precipitation, and drought were the most important drivers globally. Mansuy et al. (2019) conducted a study across the U.S. and Canada and found that climate variables were the primary control factors for the burned area in most ecoregions, both inside and outside protected areas, surpassing landscape and anthropogenic factors. The climate data used in their study included annual average precipitation, annual average relative humidity, the Hargreaves climate moisture deficit (an annual measure of energy and moisture), and degree days below 18°C. Abdollahi and Pradhan (2023) conducted a wildfire prediction study in the Gippsland region of Victoria, Australia, and found that humidity, wind speed, and rainfall were the most influential factors in wildfire prediction. Zhao and Liu (2021) used data from over 600 meteorological stations in forest, grassland, and shrubland areas to calculate two drought indices, the Keetch-Byram Drought Index and the Standardized Precipitation Index. They then used global-scale PDSI data to mask these indices and assess long-term wildfire risk in China. The results showed that drought indices had a stronger correlation with seasonal and annual fire occurrences compared to climate indicators such as temperature and precipitation.
@@ 4.4 Wildfire historical records
Wildfire historical records are indispensable in all studies and form the data foundation for wildfire risk assessment models. The sources and scales of wildfire historical records vary widely. For example, in Canada, many provinces, cities, or forestry management agencies maintain their own wildfire historical records. Additionally, the Canadian Natural Resources Department maintains its wildfire information data system, such as the Canadian Wildfire Information System. Specifically, commonly used public global or regional scale wildfire historical records include the National Burned Area Composite dataset (NBAC), the Canadian National Fire Database (CNFDB), the National Institute for Space Research database (INPE), the Fire Planning and Analysis Fire Occurrence Database (FPA-FOD), the Global Fire Emissions Database (GFED), and FireCII. In regions without systematic wildfire records, most studies use satellite-derived thermal anomalies or wildfire products from MODIS, Suomi-National Polar-orbiting Partnership (S-NPP), National Oceanic and Atmospheric Administration (NOAA) VIIRS, GOES, AVHRR, Sentinel, Landsat, Himawari, FY-4, among others.

### v4

@@ 4.1.3 Fuel types
Another method to describe fuel properties, aside from fuel moisture code and fuel load, is by categorizing fuels into homogeneous groups that exhibit similar characteristics under specified burning conditions. These groups are defined as "fuel types" or "fuel models" (Arroyo et al., 2008). The description of fuel types is generally complex, taking into account various fuel characteristics such as vegetation type, canopy height, canopy base height, canopy bulk density, vegetation cover ratio, horizontal and vertical continuity, fuel load, fuel moisture, and biomass (Chuvieco et al., 2003; Pettinari and Chuvieco, 2020; Aragoneses and Chuvieco, 2021). In recent years, various standardized fuel classification systems have been proposed in different regions (Table 1), including the NFDRS (Cohen, 1985), Northern Forest Fire Laboratory (NFFL) (Albini, 1976; Burgan, 1984), Fuel Characteristics Classification System (FCCS) (Sandberg et al., 2001; Ottmar et al., 2007), McArthur fuel types (McArthur, 1966, 1967), Canadian Fire Behaviour Prediction (FBP) (Group et al., 1992), and Prometheus (Riaño et al., 2002).
@@
4.1.4 Fuel continuity
CBH is defined as the height above which there is sufficient canopy fuel per unit volume to support vertical fire spread into the canopy (Scott, 2001). It has widespread application in various wildfire behavior models (Hall and Burke, 2006). CBH can be estimated using allometric equations based on species, diameter at breast height (DBH), and tree height, or it can be retrieved from remote sensing data. However, due to the limitations of passive optical remote sensing signals and synthetic aperture radar (SAR) in penetrating the canopy, studies using multi-angle observations or bidirectional reflectance distribution function (BRDF) retrievals to obtain three-dimensional vegetation structure information are rare in wildfire prediction (Pisek et al., 2015; Canisius and Chen, 2007). More commonly, LiDAR is used to estimate CBH at the individual tree or local scale (Andersen et al., 2005; Luo et al., 2018; Wang et al., 2022a). Some studies suggest that while the definition of CBH is clear, its practical application is complicated by the lack of a clear fuel threshold per unit volume and the potential neglect of fine fuels, which can lead to errors in modeling canopy fire risk (Cruz et al., 2004; Kramer et al., 2014). As a result, a quick and simple method to estimate ladder fuels is by estimating the coverage area or coverage rate of low-lying fuels as a representation of ladder fuel (Clark et al., 2009; Wing et al., 2012).
In summary of fuel conditions, it is noteworthy that various fuel state description parameters are generally used in studies evaluating their effectiveness in wildfire risk modeling but are rarely directly utilized in wildfire risk assessment models. Commonly used wildfire risk assessment models often directly employ raw multispectral or microwave data or indices derived from these data. This could be due to several reasons:
@@ 4.1.4 Fuel continuity
Lastly, data-driven wildfire risk prediction algorithms are becoming increasingly popular, and these algorithms can either rely on or bypass manual feature engineering, allowing them to directly learn the necessary abstractions from features or raw data.
@@ 4.2 Weather and climate conditions
Weather and climate factors are major drivers of wildfire activity, influencing the occurrence, burned area, and fire behavior (Swetnam and Betancourt, 1990; Bessie and Johnson, 1995; Abatzoglou and Kolden, 2013). Weather conditions are central to almost all global fire danger risk systems. High temperatures, low relative humidity, insufficient precipitation, and strong winds have been proven to be key factors in the ignition of fires in the short term (Keane et al., 2001; Zacharakis and Tsihrintzis, 2023b), and they also affect fire size and spread (Abatzoglou and Kolden, 2011). On the other hand, climate conditions generally refer to the long-term weather conditions of a region and their variations. Some studies specifically define these as the collective atmospheric conditions months or even years before the start of the fire season, emphasizing the use of prior climate conditions to convert fuel conditions into wildfire potential (Abatzoglou and Kolden, 2013). This approach enables predictions of large-scale fire risks over seasonal, annual, or even multi-decadal time scales.
@@ 4.2.1 Weather conditions
Short-term weather conditions focus primarily on predicting the impact of weather conditions over quarters, months, weeks, or days on fire risk. These include short-term temperature, wind, relative humidity, soil moisture, precipitation, and lightning, as well as corresponding indices like the Palmer Drought Severity Index (PDSI) and the FWI. Temperature and various humidity conditions affect fuel conditions, such as evapotranspiration and moisture content, which in turn influence the wildfire risk. Wind direction, wind speed, and other related metrics like Mean Wind Power Density (MWPD) and Mean Wind Speed (MWS) affect fire spread and intensity, making them commonly used parameters in wildfire prediction. Unlike other parameters, lightning can act as an ignition factor for wildfires.
Short-term weather parameters are widely used in wildfire risk prediction. For example, Tavakkoli Piralilou et al. (2022) used temperature, precipitation, MWPD, and MWS as parameters in their wildfire prediction model for the forested areas of Gilan Province, Iran. In another study conducted in Pakistan, Kanwal et al. (2023) utilized 16 meteorological parameters, including monthly average precipitation, evapotranspiration, wind speed, soil temperature, humidity, heat flux, albedo, average land surface temperature, soil moisture, actual evapotranspiration, moisture deficit, downward surface shortwave radiation, precipitation accumulation, minimum/maximum temperature, and vapor pressure. An assessment of their importance indicated that heat flux, evapotranspiration, and vapor pressure were relatively more important compared to other meteorological parameters, although the differences in importance were not significant. Bergado et al. (2021) used seven meteorological factors in their wildfire prediction study for Victoria, Australia: daily maximum temperature, daily minimum relative humidity, daily total solar radiation, daily total precipitation, daily average wind speed, daily average wind direction, and lightning frequency. The lightning data was derived from an annual lightning climatology dataset that provided estimates of lightning density for each day of the year. Their study found that total precipitation, lightning density, and surface temperature consistently held high weight in all models. In contrast, Quan et al. (2023) used various meteorological variables from the ERA5 dataset, including precipitation, relative humidity, temperature, wind speed, and the Keetch-Byram Drought Index, to predict wildfire risk in parts of the western Tibetan Plateau in China. The study revealed that relative humidity was the most critical factor in the occurrence of forest and grassland fires in the region. Higher relative humidity leads to higher DFMC, thereby reducing the likelihood of ignition.
@@ 4.2.2 Climate conditions
Climate conditions primarily refer to annual or multi-year meteorological parameters such as temperature, humidity, precipitation, and evaporation. Unlike short-term meteorological conditions that affect fuel status temporarily, climate conditions play a crucial role in determining fuel types, distribution, life cycles, and other factors that influence fuel characteristics, such as fuel load and continuity (Aldersley et al., 2011; Jaafari et al., 2017; Tavakkoli Piralilou et al., 2022). These fuel characteristics respond to climatic patterns accumulated before the start of the fire season (Chen et al., 2016). Consequently, climate variables are often the primary drivers of long-term wildfire activity (Pollina et al., 2013; Jain et al., 2020). Furthermore, climate conditions also explain strong seasonal variations in wildfire risk in certain regions (Martell et al., 1989), particularly in the context of intensified climate change in recent years, which has raised concerns about its long-term impact on the occurrence of large-scale wildfires (Prior and Eriksen, 2013; Jolly et al., 2015; Jain et al., 2024).
@@
Climate conditions are widely used in wildfire risk prediction. Specifically, Chen et al. (2016) used large-scale sea surface temperature anomalies to predict the severity of the fire season in South America on a seasonal time scale. Their study found that using a single ocean climate index could predict about 48% of the burned area globally for up to three months or longer before the peak burning month. In a study by Jaafari et al. (2019) in the Zagros ecoregion of western Iran, they used three climate conditions: annual temperature, rainfall, and wind effects. However, the results showed that only the annual temperature parameter was relatively important. Similarly, Naderpour et al. (2020) studied the northern beaches of New South Wales, Australia, using three climate factors: annual rainfall, annual humidity, and annual wind speed. Holdrege et al. (2024) studied the Great Basin sagebrush regions across 13 western U.S. states and used three climate factors: annual mean temperature, annual precipitation, and the proportion of summer (June to August) precipitation to total annual precipitation. The study found that areas with low summer precipitation proportion, medium to high annual precipitation, and high temperatures had the highest wildfire probability. Similarly, Aldersley et al. (2011) used annual cumulative precipitation, annual lightning frequency, monthly precipitation frequency, monthly average temperature, and climate data from the British Atmospheric Data Centre (BADC) to study the drivers of monthly fire area on a global and regional scale. They found that climate factors such as high temperatures, moderate precipitation, and drought were the most important drivers globally. Mansuy et al. (2019) conducted a study across the U.S. and Canada and found that climate variables were the primary control factors for the burned area in most ecoregions, both inside and outside protected areas, surpassing landscape and anthropogenic factors. The climate data used in their study included annual average precipitation, annual average relative humidity, the Hargreaves climate moisture deficit (an annual measure of energy and moisture), and degree days below 18°C. Abdollahi and Pradhan (2023) conducted a wildfire prediction study in the Gippsland region of Victoria, Australia, and found that humidity, wind speed, and rainfall were the most influential factors in wildfire prediction. Zhao and Liu (2021) used data from over 600 meteorological stations in forest, grassland, and shrubland areas to calculate two drought indices, the Keetch-Byram Drought Index and the Standardized Precipitation Index. They then used global-scale PDSI data to mask these indices and assess long-term wildfire risk in China. The results showed that drought indices had a stronger correlation with seasonal and annual fire occurrences compared to climate indicators such as temperature and precipitation.
@@ 4.3 Socio-economic factors
4.4 Terrain and hydrological features
Terrain factors used in wildfire prediction studies include elevation, slope, aspect, hill shade, and planar curvature. Terrain indirectly determines mesoscale and microscale climate and airflow (Coen et al., 2013), thereby impacting fuel conditions and wildfire propagation. For example, to some extent, an increase in slope leads to faster fire spread (Ghorbanzadeh and Blaschke, 2018). Steeper slopes result in faster fuel preheating and ignition rates, and compared to flat terrain, the rate of fire spread on a 20° slope increases by about four times (Lecina-Diaz et al., 2014). The slope aspect influences prevailing winds, humidity, solar radiation, and plant species distribution. In many regions of the Northern Hemisphere, north-facing slopes are typically cooler and more moist than south-facing slopes, making south-facing slopes at higher risk of forest fires (Sayad et al., 2019). Curvature and valley depth are geomorphic indicators of soil moisture and vegetation distribution (Kalantar et al., 2020). In larger valleys, downslope-driven cold air pooling has a significant impact on vegetation density and composition (Kiefer and Zhong, 2015). Additionally, elevation can also influence human activities, as lower elevations tend to have higher human population density and are more prone to wildfires compared to higher elevation areas (Guo et al., 2016). In addition, the main hydrological feature is the distance between fuel and water bodies, soil moisture, and the topographic wetness index (TWI). TWI represents soil saturation and surface runoff rates. Land with higher TWI is more saturated, and high moisture content can prevent fire ignition (Achu et al., 2021).
For example, Ghorbanzadeh et al. (2019) studied the forestry area of Amol County in the Mazandaran province of northern Iran and used slope, slope aspect, and altitude as terrain factors. They found that steep areas were generally more prone to forest fires. Nami et al. (2018) considered terrain factors such as slope, slope aspect, altitude, planar curvature, TWI, and topographic roughness index (TRI) when mapping wildfire risk in the Hyrcanian ecological region in northern Iran. Their further collinearity analysis revealed significant collinearity between TWI and slope with other parameters, leading to their exclusion from subsequent modeling tasks. A study conducted by Tran et al. (2023) on the island of Oahu in the Hawaiian Islands also utilized terrain and hydrological features such as elevation, slope, aspect, plan curvature, profile curvature, valley depth, TWI, and proximity to rivers. The assessment of their significance indicates that slope, elevation, and TWI are of secondary importance compared to proximity to roads and temperature in relation to fire risk. Gentilucci et al. (2024) also used TWI in a study in the Marche region of central Italy. Their research shows that areas of the terrain with less accumulation of moisture are most susceptible to fires.
Soil moisture information can serve as a complement or substitute for drought indices and is commonly used in wildfire risk prediction models. It can be utilized in modeling fuel moisture and other related factors (Krueger et al., 2022). For example, Bakke et al. (2023) conducted a study in Fennoscandia, Finland, to explore the primary hydro-meteorological factors influencing wildfire occurrence. They found that shallow soil moisture anomalies were the main influencing factor. Similarly, in the study by Vissio et al. (2023) on the prediction of summer burning area in Italy, the ERA5 soil moisture reanalysis data product was used. The research found that using only a single variable, namely surface soil moisture, for predicting wildfire burn area yielded good results. This is because surface soil moisture can serve as a proxy for the combined effects of temperature and precipitation. In the study by Nur et al. (2023) focusing on Sydney, Australia, the MODIS Terra Climate dataset was used, which provided drought indices and soil moisture data. However, due to the low resolution of the dataset (4 km), the importance of soil moisture as a factor was not fully reflected. In research conducted by Yahia et al. (2023), to improve the resolution of soil moisture, temperature vegetation dryness index (TVDI), perpendicular drought index (PDI), optical trapezoid model (OPTRM), and modified normalized difference water index (MNDWI) were used as proxies for soil moisture. These variables, along with a Gaussian Naive Bayes (GNB) classifier, were employed to generate wildfire risk maps. Beyond, the NDVI can be used to estimate soil moisture anomalies (Tucker, 1979; Liu et al., 2010). However, when modeling moisture anomalies using NDVI, it can only directly detect surface soil moisture, and a physical or empirical relationship must be employed to further obtain deep soil moisture. In addition, it is worth noting that the estimation of soil moisture based on optical or SAR remote sensing imagery still faces significant challenges. This is due to the influence of factors such as tree canopy obstruction, absorption, reflection, and scattering, which result in a mixed signal received by the sensor containing information on both soil and fuel moisture (Pelletier et al., 2023).
@@ 4.5 Wildfire historical records
Wildfire historical records are indispensable in all studies and form the data foundation for wildfire risk prediction models. The sources and scales of wildfire historical records vary widely. For example, in Canada, many provinces, cities, or forestry management agencies maintain their own wildfire historical records. Additionally, the Canadian Natural Resources Department maintains its wildfire information data system, such as the Canadian Wildfire Information System. Specifically, commonly used public global or regional scale wildfire historical records include the National Burned Area Composite dataset (NBAC), the Canadian National Fire Database (CNFDB), the National Institute for Space Research database (INPE), the Fire Planning and Analysis Fire Occurrence Database (FPA-FOD), the Global Fire Emissions Database (GFED), and FireCII. In regions without systematic wildfire records, most studies use satellite-derived thermal anomalies or wildfire products from MODIS, Suomi-National Polar-orbiting Partnership (S-NPP), National Oceanic and Atmospheric Administration (NOAA) VIIRS, GOES, AVHRR, Sentinel, Landsat, Himawari, FY-4, among others.
@@
Figure 3: Feature utilization frequencies is classified into four categories: fuel conditions, weather and climate conditions, socio-economic variables, and terrain and hydrological attributes. The color scheme in this visual representation signifies the prevalence of these features, with red, green, and blue denoting utilization frequencies exceeding 20, 10, and 0 occurrences, respectively.